myininaya (myininaya):

hey someone asked this question earlier, I would like to know how to solve it what is the exact values of y(2) if y'=sin(xy)

myininaya (myininaya):

i been stuck on it for the longest time

OpenStudy (amistre64):

what does y = cos(xy) derive to?

OpenStudy (amistre64):

not that its the answer

OpenStudy (amistre64):

y' = (x'y + xy')-sin(xy) right?

OpenStudy (amistre64):

y + x(sin(xy) = 1then maybe

myininaya (myininaya):

y'=(x'y+xy')(-sin(xy))

OpenStudy (amistre64):

lets assune it was -c0s :)

OpenStudy (amistre64):

then y + x(sin(xy)) = 1 would be good right?

OpenStudy (amistre64):

y = -sin(xy) x +1 ..... dunno if that helps but it might :)

myininaya (myininaya):

he said: the final answer is suppose to be y(2)=1.632 lets see if that equation holds for his answer

myininaya (myininaya):

do you know anything about euler's method?

OpenStudy (amistre64):

only by name; havent rad to much into it tho

myininaya (myininaya):

i don't i can do this problem right now my head hurts from trying to find y(2). i

myininaya (myininaya):

thanks amistre

OpenStudy (amistre64):

gets some vittles :)